The area of a rectangle is 48 square centimeters. The length of the rectangle is 8 centimeters longer than the width.

Which equation models this situation?


8w = 48


w + 8w = 48


w(w+8)=48

w + 8 = 48
1
Expert's answer
2020-10-05T16:38:27-0400


Area of rectangle is a*b=length *width,

A=48cm2,,

a=w cm

b=w+8cm because it is longer.

Therefore A= w(w+8)=48

Answer: w(w+8)=48.